How many moles of ATP can be gained from the catabolism of the following substrates to pyruvate:
a. 4 Mole mannose
b. 1 Mole lactose
c. 5 mole glycerol
1
Expert's answer
2020-11-12T06:29:43-0500

a. 5*64(60)=320(300)ATP

b. 64 (60) ATP

c. 2(4)ATP*5=10(20)ATP